Find the general/particular solution of the following Differential Equations


(Non-Exact D.E)


(2x² - 2y² + 2xy)dx + (x² - 2y)dy=0


1
Expert's answer
2021-11-15T18:15:04-0500

"M=2x^2-2y^2+2xy"

"\frac{dM}{dy}=-4y+2x=2x-4y"

"N=x^2-2y"

"\frac{dN}{dx}=2x"

Clearly;

"\frac{dM}{dy}\neq\frac{dN}{dx}"

The equation is not exact.

Use "e^{2x}" as the integrating factor to make the equation exact.

The solution will be ;

"U(x,y)=\int e^{2x}(x^2-2y)dy"

"U(x,y)=e^{2x}x^2\int1dy-2e^{2x}\int ydy"

"U(x,y)=x^2e^{2x}y-e^{2x}y^2+C"